OpenStudy (anonymous):

The following problem refers to strings in A, B, ..., Z. How many four-letter strings are there that contain exactly two J's? (Hint: Consider the disjoint cases determined by where the J's are in the string.)

OpenStudy (anonymous):

4C2*26*26=6*26*26

OpenStudy (anonymous):

4C2 is the number of ways to place the 2 Js, 26*26 is the number of ways to place the other two letters

OpenStudy (anonymous):

I'm a little confused, I keep getting 745 on my calculator, please explain if possible

OpenStudy (anonymous):

6*26*26=4056, not sure why your calculator is giving you 745

OpenStudy (anonymous):

Okay, I kept putting in 4 instead of 6 thx

OpenStudy (anonymous):

BTW I just tried 4056, my assignment is telling me its not the correct answer, I feel that it should have been 25*25 because of the 2 J's

OpenStudy (anonymous):

actually you're right. it's 6*25*25

OpenStudy (anonymous):

Yes, got it, thk you so much (high five)